# LIFBench: Evaluating the Instruction Following Performance and Stability of Large Language Models in Long-Context Scenarios

### v2

@@ 5.3 Effects of longer inputs
The negative impact of input length is also evident in stability metrics. As illustrated in Figure 4(b), most models demonstrate poorer stability in long-context scenarios, with the lowest stability observed at the longest input lengths. Interestingly, some models, such as C4AI-cmd-r-08-2024 and GLM-4-9b-chat-1m, exhibit the least stability at mid-range input lengths (e.g., 64k tokens), diverging from their overall performance trends. Additionally, the sensitivity to input length also varies across different perspectives. For instance, models like C4AI-cmd-r-08-2024 and Qwen2.5-7B-Inst. show more significant upward trend in "Expression" compared to "Variable". This highlights potential areas for improvement in enhancing instruction-following stability under long input contexts.
@@
Model size also plays a crucial role in long-context instruction-following performance. Larger models generally are better at capabilities like "Ori," "Num," and "Spat," indicating that handling original content and spatial reasoning demands more model parameters. In contrast, format-related abilities are effectively managed even by smaller models, suggesting they are less dependent on model size.
@@ Appendix B Effectiveness of Expression Extension
Implementation details   To implement RECS for expression extension, we chose the BGE-M3 text embedding model Chen et al. (2024) for encoding and applied K-means clustering Macqueen (1967). Additionally, since LLMs can introduce inaccuracies in rewrites due to misinterpretations, we manually filtered out unsuitable rewrites during the sampling phase, resulting in 5~10 instruction templates for each task.
@@
As shown in 6, compared to random sampling, RECS led to higher expression IFS scores in 75% of the models, outperforming the IFS improvement rates observed for the Variable and Length perspectives, at 37.5% and 62.5% respectively. Additionally, when comparing the mean IFS values across all models, we observed that RECS improved expression IFS while slightly lowering variable IFS and length IFS. This suggests that the RECS method make it more challenging for models to maintain stability when handling varied expressions, which shows the effectiveness in expression extension as well.
@@ Appendix C Sampling Space for Instruction Variables
As shown in Figure 8, we reserved identical placeholders for each instruction template and set the sampling space accordingly. Depending on the task requirements, the types of instruction variables include numerical values, lists, phrases, long sentences, and format indicators, etc.
@@
Appendix D Rubric Design and Program Evaluation
Rubric Design   As shown in Table 5 and Figure 2, we manually designed a scoring rubric for LIFBench, ensuring a balanced evaluation across different scenarios and capabilities. The rubric assigns greater weight $\tilde{s}$ to score point $s$ that is more challenging. For instance, in assessing format correctness, compared to the QA (OQ) task, the Batch-Label (MB) task are assigned higher scoring weights due to more complex formatting requirements.
@@ Appendix D Rubric Design and Program Evaluation
Program Evaluation   As described in Section 4.1, we design an automated evaluation program based on the scoring rubric to evaluate the quality of LLM responses. Inspired by Zhou et al. (2023a), during task design, we aim to ensure that the correct answers could be captured by the program. However, our tasks incorporate more complex formatting and logical constraints, making it challenging to directly obtain accurate evaluation results through simple automated methods. To address this, we iteratively decomposed and refined each scoring point $s$ into sub-evaluation criteria that could be directly assessed by the program, enabling a more detailed and discriminative evaluation.
@@
Our program evaluation adheres to two principles: (1) correct answers should achieve full scores, and (2) partially correct answers should be distinguishable in terms of scores. In general, we first extract structured information from the outputs based on the format constraints of each task, and then perform evaluations on other dimensions using this information. However, LLM responses are not always well-structured and may contain minor errors that hinder the extraction of structured information. To mitigate this, we carefully designed task-specific validation programs that employ techniques such as regex matching and substring retrieval to maximize the extraction of valid part from LLM responses. Finally, to ensure evaluation quality, we manually provided reference answers for all examples in LIFBench, verifying that they can achieve full scores in our automated evaluation process.
Appendix E Inspiration for IFS
@@
Appendix F Experiments Setup
Baselines   For closed-source models, we specifically use gpt-4o-2024-08-06 and gpt-4-0125-preview to represent the performance of GPT-4o and GPT-4, respectively. For open-source models, we test the performance of both base models and fine-tuned models. Models with the suffix "-Inst." or "-chat" indicate that they have been fine-tuned on instruction or dialogue data.
Inference   During the inference process, we complete the deployment of all open-source models with the vLLM Kwon et al. (2023). The temperature is set to 0 to ensure deterministic outputs. For the SFT versions of the models, we used the official chat template. For the base versions, we add a suffix "Output: " to prompt the models to generate answers according to the instructions. In order to avoid unnecessary time consumption resulting from endless repetitions in the output, we set maximum generation lengths for different tasks: List scenario tasks that output a single element were limited to 100 tokens, tasks in the MultiDoc scenario are set to 4096 tokens, and the rest of the tasks are limited to 512 tokens.
Due to variations in encoding efficiency across models’ tokenizer, some models can not accommodate the longest inputs. To address this, we apply right truncation to the context $X$, ensuring the completeness of scenario description $D$ and instruction $I$.
@@ Appendix F Experiments Setup
Number of output elements is correct. 3 Num

### v3

@@ 5.3 Effects of longer inputs
The negative impact of input length is evident in stability metrics. As illustrated in Figure 4(b), most models demonstrate poorer stability in long-context scenarios, with the lowest stability observed at the longest input lengths. Interestingly, some models, such as C4AI-cmd-r-08-2024 and GLM-4-9b-chat-1m, exhibit the least stability at mid-range input lengths (e.g., 64k tokens), diverging from their overall performance trends. Additionally, the sensitivity to input length also varies across different perspectives. For instance, models like C4AI-cmd-r-08-2024 and Qwen2.5-7B-Inst. show more significant upward trend in "Expression" compared to "Variable". This highlights potential areas for improvement in enhancing instruction-following stability under long input contexts.
@@
Model size also plays a crucial role in long-context instruction-following. Larger models generally are better at capabilities like "Ori," "Num," and "Spat," indicating that handling original content and spatial reasoning demands more model parameters. In contrast, format-related abilities are effectively managed even by smaller models, suggesting they are less dependent on model size.
5.4 Impact of Instruction Complexity
To analyze how context length and instruction complexity simultaneously affect model performance, we divide tasks into two groups based on scoring weight: easy ($\leq$10) and hard (>10). As noted in Section 4.1, higher weights indicate greater challenges, with 10 chosen to balance the number of tasks in each group.
As shown in Figure 6, the performance of different models varies significantly. Llama-3.1-70B-Inst. struggles with long contexts across all tasks, highlighting its difficulty in handling long inputs. For GPT-4 and Qwen2.5-72B-Inst., the performance on both groups is similar in short contexts, but has sharper declines on hard tasks with longer contexts, revealing their limitations in handling complex tasks in long-context scenarios. Other models exhibit larger performance gaps even in short input (i.e., 4k tokens), indicating greater task difficulty impact for them. Generally, hard group degrades more significantly as context length increases, suggesting a compounded negative effect of task complexity and context length on LLM performance.
@@ 6 Conclusion
Acknowledgments
This work was supported in part by National Natural Science Foundation of China (No. 92270119) and Key Laboratory of Advanced Theory and Application in Statistics and Data Science, Ministry of Education.
@@ Appendix B Effectiveness of Expression Extension
Implementation details   To implement RECS for expression extension, we chose the BGE-M3 text embedding model Chen et al. (2024a) for encoding and applied K-means clustering Macqueen (1967). Additionally, since LLMs can introduce inaccuracies in rewrites due to misinterpretations, we manually filtered out unsuitable rewrites during the sampling phase, resulting in 5~6 instruction templates for each task.
@@
As shown in 7, compared to random sampling, RECS led to higher expression IFS scores in 75% of the models, outperforming the IFS improvement rates observed for the Variable and Length perspectives, at 37.5% and 62.5% respectively. Additionally, when comparing the mean IFS values across all models, we observed that RECS improved expression IFS while slightly lowering variable IFS and length IFS. This suggests that the RECS method make it more challenging for models to maintain stability when handling varied expressions, which shows the effectiveness in expression extension as well.
@@ Appendix C Sampling Space for Instruction Variables
As shown in Figure 9, we reserve identical placeholders for each instruction template and set the sampling space accordingly. Depending on the task requirements, the types of instruction variables include numerical values, lists, phrases, long sentences, and format indicators, etc. We carefully consider variable distribution to avoid biases from the sampling mechanism. For example, in position-related variables within the List scenario, inputs were divided into three sections (beginning, middle, and end), with the middle section receiving the largest portion. We then randomly sample 2–3 elements from each section to ensure balance. Additionally, we introduce unconventional causal relationships in certain rule-based variables to increase task difficulty. For instance, in the QA task, models are required to use "False" for correct answers and "True" for incorrect ones, making it more challenging to follow instructions.
@@
Appendix D Rubric Design and Evaluation Program
Rubric Design   For the weight $\tilde{s}$ assignment of scoring points, we primarily considered three factors:
(1) Evaluation complexity: Scoring points with higher evaluation complexity require more assessment steps and are therefore assigned greater weight. For example, in assessing format correctness, the Batch-Label (MB) task is given higher scoring weights compared to the QA (OQ) task due to its more complex formatting requirements.
(2) Task difficulty for models: Some scoring points may not be as complex to evaluate but are particularly challenging for the model to fulfill. In such cases, we allocate greater weight to these points as well.
(3) Balance across capabilities and scenarios: Based on the considerations in (1) and (2), we made fine adjustments to ensure that the weights assigned to different instruction-following capabilities and scenarios are as balanced as possible. This balance is shown in Figure 2.
On this basis, we present the rubric of LIFEval in Table 6. Additionally, the pseudocode examples reflect our considerations above. Program 1 show the correspondence between evaluation steps and their respective weights. In Program 2, we award an additional point for responses that fully meet the requirements, highlighting the difficulty of that particular scoring point.
@@ Appendix D Rubric Design and Evaluation Program
Evaluation Program   As described in Section 4.1, we design an automated evaluation program based on the scoring rubric to evaluate the quality of LLM responses. Inspired by Zhou et al. (2023a), during task design, we aim to ensure that the correct answers could be captured by the program. However, our tasks incorporate more complex formatting and logical constraints, making it challenging to directly obtain accurate evaluation results through simple automated methods. To address this, we iteratively decomposed and refined each scoring point $s$ into sub-evaluation criteria that could be directly assessed by the program, enabling a more detailed and discriminative evaluation.
@@
Our program evaluation adheres to two principles: (1) correct answers should achieve full scores, and (2) partially correct answers should be distinguishable in terms of scores. In general, we first extract structured information from the outputs based on the format constraints of each task, and then perform evaluations on other dimensions using this information. However, LLM responses are not always well-structured and may contain minor errors that hinder the extraction of structured information. To mitigate this, we carefully designed task-specific validation programs that employ techniques such as regex matching and substring retrieval to maximize the extraction of valid part from LLM responses (see Program 1). Finally, we manually provided reference answers for all samples in LIFBench. We verified that the automated evaluation programs in LIFEval consistently achieve full scores on all reference answers, thereby fulfilling the first principle outlined above.
Appendix E Experiments Setup
Baselines   For closed-source models, we specifically use gpt-4o-2024-08-06 and gpt-4-0125-preview to represent the performance of GPT-4o and GPT-4, respectively. For open-source models, all implementations were sourced from Hugging Face [^5] https://huggingface.co/ (as detailed in Table 5), and we test the performance of both base models and fine-tuned models. Models with the suffix "-Inst." or "-chat" indicate that they have been fine-tuned on instruction or dialogue data.
Inference   During the inference process, we complete the deployment of all open-source models with the vLLM Kwon et al. (2023). The temperature is set to 0 to ensure deterministic outputs. For the SFT versions of the models, we used the official chat template. For the base versions, we add a suffix "Output: " to prompt the models to generate answers according to the instructions. Token counts were calculated using GPT-4’s tokenizer [^6] https://github.com/openai/tiktoken, and truncation was applied to adjust context for models unable to process the longest contexts. To ensure diversity and accuracy in instruction phrasing, we generated 5–6 template variants for each original instruction and manually filtered them. Each template included placeholders for variables (e.g., numerical values, lists, phrases, long sentences, and format indicators), dynamically sampled from curated sets of 5–10 candidates to create task-specific prompts. In order to avoid unnecessary time consumption resulting from endless repetitions in the output, we set maximum generation lengths for different tasks: List scenario tasks that output a single element were limited to 100 tokens, tasks in the MultiDoc scenario are set to 4096 tokens, and the rest of the tasks are limited to 512 tokens.
Due to variations in encoding efficiency across models’ tokenizer, some models can not accommodate the longest inputs. To address this, we apply right truncation to the context $X$, ensuring the completeness of scenario description $D$ and instruction $I$. The experiments were conducted across six context length intervals, ranging from 4k to 128k tokens, with task-specific output limits to ensure sufficient space for model generation.
  Model Hugging Face ID
Llama-3.1-70B-Inst. meta-llama/Llama-3.1-70B-Instruct
Llama-3.1-8B-Inst. meta-llama/Llama-3.1-8B-Instruct
Llama-3.1-70B meta-llama/Llama-3.1-70B
Llama-3.1-8B meta-llama/Llama-3.1-8B
[…]
Qwen2.5-32B-Inst. Qwen/Qwen2.5-72B-Instruct
Qwen2.5-14B-Inst. Qwen/Qwen2.5-72B-Instruct
Qwen2.5-7B-Inst. Qwen/Qwen2.5-72B-Instruct
Qwen2.5-72B Qwen/Qwen2.5-72B
Qwen2.5-32B Qwen/Qwen2.5-72B
Qwen2.5-14B Qwen/Qwen2.5-72B
Qwen2.5-7B Qwen/Qwen2.5-72B
[…]
C4AI-cmd-r-v01 CohereLabs/c4ai-command-r-v01
[…]
LWM-Text-1M LargeWorldModel/LWM-Text-1M
[…]
Table 5: A mapping between the Hugging Face model IDs and the aliases used in the paper.
Appendix F Inspiration for IFS
@@
Program 1 Format Validation (JSON Dict)
Target: Verify if the output meets the JSON Dict format.
1:LLM answer $a$, full score $\tilde{s}=4$
2:Validation score $s_{a}=f_{s}(a)\in[0,\tilde{s}]$
3:Step 1: Symbol Check (Max 1 point)
4:$s_{a}\leftarrow 0$
[…]
6:     $s_{a}\leftarrow s_{a}+1$
7:end if
8:Step 2: Parsing Check (Max 2 points)
[…]
11:     $s_{a}\leftarrow s_{a}+2$
[…]
13:     $s_{a}\leftarrow s_{a}+1$
14:end if
15:Step 3: KV Check (Max 1 point)
[…]
17:     $s_{a}\leftarrow s_{a}+1$
18:end if
19:return $s_{a}$
[…]
Program 2 Quantity Verification (LMI)
Target: Verify if the number of output elements is correct.
1:LLM answer $a$, target number $n_{\text{gold}}$, full score $\tilde{s}=3$
2:Validation score $s_{a}=f_{s}(a)\in[0,\tilde{s}]$
3:$n_{\text{pred}}\leftarrow\text{extract\_quantity}(a)$
4$\triangleright$ Extract the number of elements in the output
5:if $n_{\text{pred}}=n_{\text{gold}}$ then
6:     $s_{a}\leftarrow\tilde{s}$triangleright Full score if the quantity matches
7:else
8:     $s_{a}\leftarrow\max\left(0,\left(1.0-\frac{|n_{\text{pred}}-n_{\text{gold}}|}{n_{\text{gold}}}\right)\right)\times 2$
9$\triangleright$ Penalty based on relative deviation
10:end if
11:return $s_{a}$
Note: The function $n_{\text{pred}}=\text{extract\_quantity}(a)$ can be executed in various ways. If the LLM answer $a$ passes format checks and is parsable, the target quantity is directly obtained. Otherwise, string processing techniques are employed to extract it with feedback.
@@
The number of output elements is correct. 3 Num